For this exercise, you can simulate the described conditions by changing the values in the Run Experiment tool of the Simulation. To be able to measure the effects on the gas volume, slide the Rspd bar in the Properties box to V (L) so that the volume bar is highlighted (turns yellow). In this way, volume becomes the only dependent variable. Note that, when volume is selected with Rspd, its value can no longer be directly controlled.

Suppose a piston automatically adjusts to maintain a gas at a constant pressure of 12.20 atm . For the initial conditions, consider 0.03 mol of neon at a temperature of 245.00 K . This gas occupies a volume of 0.05 L under those conditions. What volume will the gas occupy if the number of moles is increased to 0.06 mol (n2) from the initial conditions? What volume will the gas occupy if the temperature is increased to 345.00 K (T2) from the initial conditions? Remember to reset the experiment to the initial conditions before determining each new volume.

Express the volumes in liters to two decimal places separated by a comma.
Answer #1

The pressure and temperature conditions are same , on increasing the number of moles of gas from 0.03 to 0.06 , volume of gas increases .

V. \alpha n (at constant Temperature and pressure)

0.05L/V2 = 0.03/0.06 = 1/2

V2 = 0.05 L × 2 = 0.10 L

--------

V. \alpha Temperature. (at constant n and pressure )

0.05L/V3 = 245.00K/345.00K =

V3 = 0.05×345.00/245.00 = 0.0704L

V3 = 0.07L

--------

The gas would occupy volume( in L) :

0.10, 0.07 (Answer)